OpenStudy (anonymous):

Find the solutions of the equation -3x^2+x-5=0

OpenStudy (anonymous):

I got the answer \[\frac{-1\pm \sqrt{59}i}{-6}\] but I don't know how to get rid of the -6 in the denominator

OpenStudy (nali):

do u have to get rid of the the -6

OpenStudy (anonymous):

yes

sam (.sam.):

\[x\text{ = }\frac{1}{6} \left(1-i \sqrt{59}\right)\] and \[x\text{ = }\frac{1}{6} \left(1+i \sqrt{59}\right)\]

sam (.sam.):

you're right

OpenStudy (anonymous):

still don't see how you got rid of the -6

sam (.sam.):

the negative 6 go with negative 1 on top simplifying your answer \[\frac{1\pm \sqrt{59}i}{6}\] and ends with my answer \[x\text{ = }\frac{1}{6} \left(1-i \sqrt{59}\right)\] \[x\text{ = }\frac{1}{6} \left(1+i \sqrt{59}\right)\]

sam (.sam.):

exactly

OpenStudy (anonymous):

nvm I get it because a -*- will give you the positive and a +*- will give you the negative

OpenStudy (anonymous):

thanks so much!!! for your efforts it was greatly appreciated!!!